# The Gray Zone of Faithfulness: Taming Ambiguity in Unfaithfulness Detection

### v2

@@ 5.1 Baselines
Fine-tuned LLMs. The fine-tuned LLMs comprise Minicheck 7B [^4] https://huggingface.co/bespokelabs/Bespoke-MiniCheck-7B (Tang et al., 2024), Lynx 8B Instruct v1.1555https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-8B-Instruct-v1.1, Lynx 70B Instruct [^6] https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-70B-Instruct (Ravi et al., 2024), and ANAH-v2777https://huggingface.co/opencompass/anah-v2 (Gu et al., 2024), where Minicheck and Lynx models are binary classifiers, and ANAH-v2 is a three-way classifier.
Zero-shot LLMs. Zero-shot LLMs are prompted with the definitions from our taxonomy (see Appendix E for the prompt). The evaluated LLMs include GPT-5, DeepSeek-R1 (DeepSeek-AI, 2025), DeepSeek-V3, Qwen3 235B A22B (Yang et al., 2025), and QwQ-32B (Team, 2025). All models decode with temperature of 0.6, except GPT-5, whose temperature is not adjustable via API.
Others. The other baselines include LLM-Check, a model-internals-based method that uses attention weights and hidden states of LLMs (Sriramanan et al., 2024), CCP (Claim Conditioned Probability), a probability-based technique (Fadeeva et al., 2024), and InFusE, a retrieve-then-verify approach (Zhang et al., 2024a). Implementation details for all baselines are provided in Appendix D.
@@ 5.2 Evaluation Protocol
Since the label space in our benchmark differs from previous work, where hallucination detection is typically treated as a binary or three-way classification, we propose two evaluation protocols to overcome this difficulty. Following Seo et al. (2025), the first protocol removes Not Sure and No-Fact instances and evaluates detectors on the remaining data, with all classes merged into two categories: Faithful and Unfaithful. This merging involves two aspects: merging annotations and merging predictions. For annotations, Explicitly-Supported and Implicitly-Supported are merged into the Faithful class, while Fabricated and Contradicting form the Unfaithful class. Predictions are then aligned accordingly: for NLI-style detectors, Entailment is mapped to Faithful, and Neutral and Contradictory to Unfaithful; for zero-shot LLMs, classes with a faithfulness degree not less than a threshold (defined later) are considered Faithful, with the rest as Unfaithful. Here, the order of faithfulness degree is defined as: $\texttt{Contradicting}\prec\texttt{Fabricated}\prec\texttt{Ambiguous}\prec\texttt{No-Fact}\prec\texttt{Out-Dependent}\prec\texttt{Implicitly-Supported}\prec\texttt{Explicitly-Supported}$. The threshold is currently set to the Implicitly-Supported class, with alternative thresholds explored in the selective prediction part of Section 5.3. After merging, we report balanced accuracy, along with hallucination detection recall, precision, and F1 (i.e., the recall, precision, and F1 for the Unfaithful class).
To fully utilize the fine-grained annotations, we propose another protocol that evaluates the ranking quality of the hallucination detectors. Without filtering or merging labels, we consider the full dataset be $\{(x_{i},y_{i})\}_{i=1}^{n}$, where $y_{i}$ is the fine-grained label of instance $x_{i}$. For predictions, instead, the output space is merged into Faithful/Unfaithful as before, assigned with a faithfulness degree order of Unfaithful $\prec$ Faithful. Let $\mathcal{P}:=\{(x_{i},x_{j})|y_{i}\succ y_{j}\}$. Inspired by the ranking expression of AUC (Area Under the Curve; Eq. (2.21) in Zhou (2021)), we propose a novel metric ranking loss:
@@
In addition, for zero-shot LLMs, we evaluate the class-wise F1 to investigate the fine-grained unfaithfulness classification performances.
Figure 4: The class-wise F1 of zero-shot LLMs.
@@ 5.3 Results
GPT-5 75.4 58.1 59.2 58.6 35.7
DeepSeek-R1 78.6 64.0 63.8 63.9 34.4
DeepSeek-V3 59.3 19.0 87.3 31.3 44.4
Qwen3 235B A22B 67.4 36.7 78.5 50.0 40.5
QwQ 32B 70.6 43.6 77.8 55.9 38.8
@@
ANAH-v2 52.8 11.1 26.9 15.7 48.8
MiniCheck 7B 61.8 27.7 55.2 36.9 42.7
Lynx 8B Instruct v1.1 62.6 38.7 34.6 36.5 42.1
Lynx 70B Instruct 67.1 40.5 54.2 46.3 41.9
@@
LLM-Check (Attn Score) 54.9 44.3 19.2 26.8 47.9
LLM-Check (Hidden Score) 55.0 56.1 18.4 27.7 48.5
InFusE 46.8 37.7 13.7 20.1 51.7
CCP 58.1 46.7 22.1 30.0 44.2
@@
The evaluation results for metrics except class-wise F1 are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, specifically GPT-5 and DeepSeek-R1, which achieve balanced accuracy of $75\%\sim 80\%$, hallucination F1 of $\sim 60\%$, and ranking loss of $\sim 30\%$. Figure 4 shows the class-wise F1 of zero-shot LLMs. The results reveal that DeepSeek-R1’s better hallucination F1 compared to GPT-5 (shown in Table 5.3) can be attributed to its higher F1 score in the Fabricated class, as this class accounts for the majority of the Unfaithful classes. Across all such models, F1 values for Ambiguous and Out-Dependent remain consistently low, suggesting that LLMs have difficulty distinguishing Not Sure instances from others. We hypothesize that for Out-Dependent, this may be due to the models’ lack of access to external knowledge sources, while their parametric knowledge is insufficient to verify such cases; and for Ambiguous, this may stem from the models’ limited sensitivity to linguistic ambiguity.
Selective Prediction Evaluation. To investigate whether the Explicitly-Supported prediction serves as a reliable indicator of faithfulness confidence, we frame zero-shot LLM detectors as confidence estimators and assess their performance under the framework of selective prediction (Geifman and El-Yaniv, 2017). In this setting, selective prediction assesses both the selective risk – the hallucination rate for samples whose confidence is greater than or equal to a certain threshold – and the coverage, defined as the proportion of samples with confidence greater than or equal to that threshold. Here, confidence is defined as the faithfulness degree, whose ordering has been established earlier. We vary the confidence threshold across Out-Dependent, Implicitly-Supported, and Explicitly-Supported to comprehensively evaluate the selective prediction performance. For a fixed coverage, a lower selective risk indicates better selective prediction performance.
The selective risk plots against coverage are shown in Figure 5, indicating that as the threshold shifts from Out-Dependent to Explicitly-Supported, selective risk decreases consistently, although coverage also drops considerably. The results highlight the potential of selective prediction in reducing risk for unfaithfulness detection, despite remaining some distance from practical application (selective risk remains > 5%).
@@ 6 Conclusion
Due to the limitation of human labor, we only consider the faithfulness over summarization tasks in our benchmark. One direction of future work of our paper could be building benchmarks with diverse knowledge-grounded tasks, such as RAG and data-to-text. Another direction of future work might be to incorporate the unfaithfulness detector with an external knowledge source, which could help improve the detection of Out-Dependent examples.
@@ Appendix A Case Study of Annotation Errors in FaithBench
Document Response Original Annotation Our Annotation
…three British writers, two Americans and an Australian on this year’s shortlist. …This year marks is 46th year of the Booker Prize. … Here’s a concise summary of the passage: The passage discusses the 2014 Man Booker Prize shortlist and its contenders. … All annotators agreed to Consistent (it seems to be Benign instead) Out-Dependent. Reason: It requires external world knowledge to tell in which year the 46th Booker Prize was awarded.
…Turkey insists that the jet, …ignored warnings to leave its airspace. Russia says it was shot down within Syrian airspace and President Vladimir Putin … …Turkey claims the jet ignored warnings to leave Syrian airspace, while Russia asserts it was shot down within Syrian airspace. … All annotators agreed to Unwanted (it seems to be Questionable instead) Contradicting / Explicitly-Supported. Reason: For ambiguous examples, we retain diverse annotations of all interpretations.
Table 5: Case study of FaithBench annotation errors, where its annotation rules of Benign and Questionable were not adhered to. Considering the original annotations are span-level, we aggregate all span-level original annotations whose span is in the target sentence to be compared with our annotations.
@@ Appendix B Details of Generating the Summaries
Appendix C Full Results of Transition Probabilities
The model-wise estimated transition probabilities and the estimated transition probabilities of every 2, 3, 4 steps are shown in Figure 6. The results are similar to the overall results in Figure 3.
@@ Appendix D Implementation Details of Baselines
The NLI model used in InFusE is the cross-encoder/nli-deberta-v3-large [^8] https://huggingface.co/cross-encoder/nli-deberta-v3-large model on HuggingFace. The CCP originally outputs an uncertainty score in [1,0][-1,0] for the LLM generation. We map the score to a 0-1 faithfulness prediction using a threshold of $\tau=-0.001$, with scores greater than $\tau$ mapped to the Unfaithful class. The LLM-Check is implemented using a base model of Llama-2-7B and a threshold of -174.0/6.3 for Attn Score / Hidden Score, respectively, where the Attn Score / Hidden Score is extracted from Layer 21/20.
Appendix E Zero-shot Detector Prompt
@@
The prompt for zero-shot LLM detectors is as follows, where <DOCUMENT>, <SUMMARY>, <SENTENCE> are the placeholders of the document, summary, and the sentence, respectively.
@@
Detector Prompt You are judging the faithfulness of the assigned sentence of a summary to the source document. The faithfulness has five options: A. Explicitly-Supported: all atomic facts of the sentence appear verbatim (up to a lexical or syntactic transformation) within the document. B. Generally-Supported: the document entails the sentence, and the sentence is not explicitly-supported by the document. Note that some difference between part of the sentence and the document is allowed, but it should be restricted to the case in which the part of the sentence adopts a weaker or less certain utterance than the document. If any part of the sentence adopts a stronger or more certain utterance than the document, consider option D (Fabricated). C. Inconsistent: the sentence logically contradicts the document. D. Fabricated: the sentence does not logically contradict the document, and is neither logically implied by the document nor by external world knowledge. E. Out-Dependent: the sentence is not logically implied by the document but by the union of the document and external world knowledge. F. Ambiguous: the sentence or the document has multiple interpretations. G. No-Fact: the sentence is devoid of facts. Your task is to decompose the target sentence into atomic facts and then output the correct option. [Document] <DOCUMENT> [Summary] <SUMMARY> [Sentence in the summary] <SENTENCE> Your option:

### v3

@@ 5.1 Baselines
Fine-tuned LLMs. The fine-tuned LLMs comprise Minicheck 7B [^5] https://huggingface.co/bespokelabs/Bespoke-MiniCheck-7B (tang2024minicheck), Lynx 8B Instruct v1.1666https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-8B-Instruct-v1.1, Lynx 70B Instruct [^7] https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-70B-Instruct (ravi2024lynx), and ANAH-v2888https://huggingface.co/opencompass/anah-v2 (gu2024anahv2), where Minicheck and Lynx models are binary classifiers, and ANAH-v2 is a three-way classifier.
Zero-shot LLMs. Zero-shot LLMs are prompted with the definitions from our taxonomy (see Appendix E for the prompt). The evaluated LLMs include GPT-5, DeepSeek-R1 (guo2025deepseekr1), DeepSeek-V3, Qwen3 235B A22B (yang2025qwen3), and QwQ-32B (team2025qwq). We also included GPT-5 enhanced with RAG (lewis2020retrieval) – hereafter referred to as GPT-5 + RAG – to provide the LLM with access to external world knowledge and help verify Out-Dependent examples. The retrieval corpus was the webpages collected during annotation (see Section 4.2). For more details, please refer to Appendix E.
Others. The other baselines include two model-internals-based methods, LLM-Check (sriramanan2024llmcheck) and ReDeEP (sun2024redeep), a probability-based method CCP (Claim Conditioned Probability) (fadeeva2024factchecing), and a retrieve-then-verify approach InFusE (zhang2024finegrained). Implementation details for all baselines are provided in Appendix D.
@@ 5.2 Evaluation Protocol
Since the label space in our benchmark differs from previous work, where hallucination detection is typically treated as a binary or three-way classification, we propose two evaluation protocols to overcome this difficulty. Following seo2025verifying, the first protocol removes Not Sure and No-Fact instances and evaluates detectors on the remaining data, with all classes merged into two categories: Faithful and Unfaithful. This merging involves two aspects: merging annotations and merging predictions. For annotations, Explicitly-Supported and Implicitly-Supported are merged into the Faithful class, while Fabricated and Contradicting form the Unfaithful class. Predictions are then aligned accordingly: for NLI-style detectors, Entailment is mapped to Faithful, and Neutral and Contradictory to Unfaithful; for zero-shot LLMs, classes with a faithfulness degree not less than a threshold (defined later) are considered Faithful, with the rest as Unfaithful. Here, the order of faithfulness degree is defined as: $\texttt{Contradicting}\prec\texttt{Fabricated}\prec\texttt{Ambiguous}\prec\texttt{No-Fact}\prec\texttt{Out-Dependent}\prec\texttt{Implicitly-Supported}\prec\texttt{Explicitly-Supported}$. The threshold is currently set to the Implicitly-Supported class, with alternative thresholds explored in the selective prediction part of Section 5.3. After merging, we report balanced accuracy, along with hallucination detection recall, precision, and F1 (i.e., the recall, precision, and F1 for the Unfaithful class).
To fully utilize the fine-grained annotations, we propose another protocol that evaluates the ranking quality of the hallucination detectors. Without filtering or merging labels, we consider the full dataset be $\{(x_{i},y_{i})\}_{i=1}^{n}$, where $y_{i}$ is the fine-grained label of instance $x_{i}$. For predictions, instead, the output space is merged into Faithful/Unfaithful as before, assigned with a faithfulness degree order of Unfaithful $\prec$ Faithful. Let $\mathcal{P}:=\{(x_{i},x_{j})|y_{i}\succ y_{j}\}$. Inspired by the ranking expression of AUC (Area Under the Curve; Eq. (2.21) in zhou2021machine), we propose a novel metric ranking loss:
@@
In addition, for zero-shot LLMs, we evaluate the class-wise precision and recall to investigate the fine-grained unfaithfulness classification performances.
@@
Figure 4: The class-wise precision and recall of zero-shot LLMs. The low recall for Ambiguous may stem from the scarcity of Ambiguous examples.
@@ 5.3 Results
DeepSeek-V3 61.2 23.0 86.8 36.4 43.4
Qwen3 235B A22B 68.9 40.1 76.7 52.6 39.7
QwQ 32B 72.4 47.4 76.8 58.6 38.0
DeepSeek-R1 78.9 64.8 62.8 63.8 33.7
GPT-5 80.0 67.9 61.1 64.4 33.0
GPT-5 + RAG 81.4 69.3 66.1 67.7 32.6
@@
ANAH-v2 53.0 11.5 27.7 16.3 46.6
MiniCheck 7B 61.8 27.9 55.2 37.0 42.8
Lynx 8B Instruct v1.1 62.6 38.7 34.6 36.5 42.3
Lynx 70B Instruct 67.2 40.8 54.4 46.6 42.0
@@
LLM-Check (Attn Score) 52.2 56.8 16.7 25.8 48.0
LLM-Check (Hidden Score) 54.9 56.1 18.3 27.6 48.7
ReDeEP 57.6 67.9 19.2 30.0 43.0
InFusE 46.6 37.3 13.5 19.8 51.9
CCP 58.6 47.7 22.4 30.5 44.1
@@
The main evaluation results are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, notably GPT-5 + RAG, which achieve balanced accuracy of over 80%, hallucination F1 of nearly 68%, and ranking loss of less than 33%. Figure 4 shows the class-wise precision and recall of zero-shot LLMs. Across all models, recalls for Ambiguous and Out-Dependent remain consistently low. Nevertheless, compared to GPT-5, GPT-5 + RAG significantly boosts the precision of Out-Dependent while maintaining the recall, showing the effectiveness of using RAG. The low recall for Ambiguous may stem from both the scarcity of Ambiguous examples and the models’ lack of sensitivity to linguistic ambiguity. As for Out-Dependent, we found this class was mainly misclassified as faithful classes (see Figure 7 in Appendix F), indicating that the low recall may be due to the LLM detectors’ unawareness of using external knowledge from their parametric memory.
Selective Prediction Evaluation. To investigate whether the Explicitly-Supported prediction serves as a reliable indicator of faithfulness confidence, we frame zero-shot LLM detectors as confidence estimators and assess their performance under the framework of selective prediction (geifman2017selective). In this setting, selective prediction assesses both the selective risk – the hallucination rate for samples whose confidence is greater than or equal to a certain threshold – and the coverage, defined as the proportion of samples with confidence greater than or equal to that threshold. Here, confidence is defined as the faithfulness degree, whose ordering has been established earlier. We vary the confidence threshold across Out-Dependent, Implicitly-Supported, and Explicitly-Supported to comprehensively evaluate the selective prediction performance. For a fixed coverage, a lower selective risk indicates better selective prediction performance.
The selective risk-coverage plots are shown in Figure 4. When the threshold transits from Out-Dependent to Explicitly-Supported, selective risk decreases consistently. GPT-5 + RAG achieves the best tradeoff between selective risk and coverage, reaching below 4% with coverage over 60%. The results highlight the potential of selective prediction in enhancing trustworthiness for unfaithfulness detection.
@@ 6 Conclusion
Due to the limitation of human labor, we only consider the faithfulness over summarization tasks in our benchmark. One direction of future work of our paper could be building benchmarks with diverse knowledge-grounded tasks, such as RAG and data-to-text. Another direction of future work might be to develop methods that can identify when an LLM leverages external world knowledge from its parametric memory, which would enhance the detection of Out-Dependent examples.
@@ Appendix A Case Study of Annotation Errors in FaithBench
Document Target Sentence Original Annot. Our Annotation
…three British writers, two Americans and an Australian on this year’s shortlist. …This year marks is 46th year of the Booker Prize. … Here’s a concise summary of the passage: The passage discusses the 2014 Man Booker Prize shortlist and its contenders. All annotators agreed to Consistent (it seems to be Benign instead). Out-Dependent. Reason: It requires external world knowledge to tell in which year the 46th Booker Prize was awarded.
…Turkey insists that the jet, …ignored warnings to leave its airspace. Russia says it was shot down within Syrian airspace and President Vladimir Putin … Turkey claims the jet ignored warnings to leave Syrian airspace, while Russia asserts it was shot down within Syrian airspace. All annotators agreed to Unwanted (it seems to be Questionable instead). Ambiguous (Contradicting / Explicitly-Supported). Reason: The annotation depends on how to interpret “its” in the document.
the plant was hit by another explosion in july 2013, although there were no reports of casualties or toxic leaks at the time . … There was no mention of toxic chemical spills in this incident, but the plant has a history of explosions, with the last one occurring in July 2013. Two of three annotators labeled the sentence as Questionable. Out-Dependent. Reason: It requires external knowledge to tell whether the explosion in July 2013 is the most recent one.
…but the veteran saw his achievement eclipsed by that of camilo villegas on a day which saw five hole-in-ones , levelling the record from 2002. the colombian followed nicklaus with an ace on the fourth and also managed one on the eighth , but he could n’t quite land the title … Despite his achievement, Camilo Villegas stole the show by scoring two holes-in-one, tying the record. One of four annotators labeled this sentence as Benign. Two of four annotators labeled this sentence as Questionable. Fabricated. Reason: The record was not the two aces of Villegas but the five aces on that day.
Table 5: Bad cases of FaithBench annotation related to Benign and Questionable classes, where Original Annot. denotes the original annotations. Considering the original annotations are span-level, we aggregate all span-level original annotations whose span is in the target sentence to be compared with our annotations.
@@ Appendix B Details of Generating the Summaries
Appendix C Sequential Dependency of Unfaithfulness
The phenomenon of hallucination snowballing (zhang2024language) – where earlier incorrect claims trigger subsequent hallucinated explanations – has been observed in several multi-step question-answering datasets, indicating a sequential dependency of hallucinations. To examine whether such sequential dependency exists in our benchmark, we model sentence-wise faithfulness in each response as a Markov chain and estimate the transition probabilities among unfaithfulness classes (see Figure 5). As the results show, the confidence intervals of $\hat{p}(\text{Faithful}\mid\cdot)$ (and similarly for $\hat{p}(\text{Unfaithful}\mid\cdot)$) overlap across all source classes (except the outlier of the No-Fact class due to data scarcity). Comparable results are observed for model-specific and longer-dependency transition probabilities. The model-wise estimated transition probabilities and the estimated transition probabilities of every 2, 3, 4 steps are shown in Figure 6. The results are similar to the overall results in Figure 5. Thus, the sequential dependency of unfaithfulness in our dataset is weak, suggesting that prior sentence hallucinations do not provide a reliable shortcut for predicting unfaithfulness in subsequent sentences.
@@ Appendix D Implementation Details of Baselines
The NLI model used in InFusE is the cross-encoder/nli-deberta-v3-large [^9] https://huggingface.co/cross-encoder/nli-deberta-v3-large model on HuggingFace. The CCP, LLM-Check, and ReDeEP originally output uncertainty scores rather than a 0-1 decision. We map the score to a 0-1 faithfulness prediction using a threshold $\tau$, with scores greater than $\tau$ mapped to the Unfaithful class. The CCP used a threshold of -0.001. The LLM-Check was implemented using a base model of Llama-2-7B (touvron2023llama) and a threshold of -174.0/6.3 for Attn Score / Hidden Score, respectively, where the Attn Score / Hidden Score is extracted from Layer 21/20. The ReDeEP was implemented with a base model of Llama-2-7B, $\alpha=1,\beta=1.6$, and a threshold of 0.6, where the selected copy heads were the top-7 scoring copy heads, and the selected FFN layers were the top-3 layers, following the settings of ReDeEP (chunk) in sun2024redeep.
Appendix E Zero-shot Detector Details
Temperature. All models decode with temperature of 0.6, except GPT-5, whose temperature is not adjustable via API.
@@ Appendix E Zero-shot Detector Details
Zero-shot Prompt. The prompt for zero-shot LLM detectors without RAG is as follows, where <DOCUMENT>, <SUMMARY>, and <SENTENCE> are the placeholders of the document, summary, and the sentence, respectively.
@@
Detector w/o RAG Prompt You are judging the faithfulness of the assigned sentence of a summary to the source document. The faithfulness has five options: A. Explicitly-Supported: all atomic facts of the sentence appear verbatim (up to a lexical or syntactic transformation) within the document. B. Generally-Supported: the document entails the sentence, and the sentence is not explicitly-supported by the document. Note that some difference between part of the sentence and the document is allowed, but it should be restricted to the case in which the part of the sentence adopts a weaker or less certain utterance than the document. If any part of the sentence adopts a stronger or more certain utterance than the document, consider option D (Fabricated). C. Inconsistent: the sentence logically contradicts the document. D. Fabricated: the sentence does not logically contradict the document, and is neither logically implied by the document nor by external world knowledge. E. Out-Dependent: the sentence is not logically implied by the document but by the union of the document and external world knowledge. F. Ambiguous: the sentence or the document has multiple interpretations. G. No-Fact: the sentence is devoid of facts. Your task is to decompose the target sentence into atomic facts and then output the correct option. [Document] <DOCUMENT> [Summary] <SUMMARY> [Sentence in the summary] <SENTENCE> Your option:
The prompt for zero-shot LLM detectors enhanced with RAG is as follows, where <DOCUMENT>, <RETRIEVED RESULTS>, <SUMMARY>, and <SENTENCE> are the placeholders of the document, retrieved results, summary, and the sentence, respectively.
Detector w/ RAG Prompt You are judging whether a given sentence extracted from a summary is faithful to its source document. Several text snippets are retrieved to provide external world knowledge of the source document. The faithfulness has five options: A. Explicitly-Supported: all atomic facts of the sentence appear verbatim (up to a lexical or syntactic transformation) within the source document. B. Generally-Supported: the source document entails the sentence, and the sentence is not explicitly-supported by the document. Note that some difference between part of the sentence and the document is allowed, but it should be restricted to the case in which the part of the sentence adopts a weaker or less certain utterance than the document. If any part of the sentence adopts a stronger or more certain utterance than the document, consider option D (Fabricated). C. Inconsistent: the sentence logically contradicts the source document. D. Fabricated: the sentence does not logically contradict the source document, and is neither logically implied by the source document nor by the retrieved snippets. E. Out-Dependent: the sentence is not logically implied by the source document but by the union of the source document and the retrieved snippets. F. Ambiguous: the sentence or the source document has multiple interpretations, which could lead to different conclusions. G. No-Fact: the sentence is devoid of facts. Your task is to decompose the target sentence into atomic facts and then output the correct option. [Source Documents] <DOCUMENT> [Retrieved Snippets] <RETRIEVED RESULTS> [Summary] <SUMMARY> [Sentence in the Summary] <SENTENCE> Your option:
RAG Details. We use the web pages collected for the Out-Dependent class during the annotation process as the retrieval corpus. The web page texts (in HTML) are first converted into Markdown, and then truncated into snippets of 1600 characters each, with overlaps of 200 characters, resulting in 3030 snippets. The query for the retrieval is the sentence to verify. For each query, we retrieved the top 3 documents using the BM25 retriever (robertson1995okapi).
Appendix F The Confusion Matrix of GPT-5 + RAG
Figure 7: The confusion matrix (with row-wise normalization) of GPT-5 + RAG.
The confusion matrix of GPT-5 + RAG is shown in Figure 7.